OpenStudy (anonymous):

if susan subtracts 11 from one- fourth of her number she gets 11. Whats susans number

OpenStudy (anonymous):

Translate your question into an equation

OpenStudy (hoblos):

x/4 -11 = 11 x/4 = 22 x=88

OpenStudy (anonymous):

x/4 - 11 = 11 x/4 = 22 x = 88

OpenStudy (anonymous):

i can't I don't know how

OpenStudy (anonymous):

So "one fourth of her number" basically means 1/4 of an unknown number, x. So 1/4x. then susan subtracts 11. So its 1/4x - 11 = 11 (because she gets 11)